Derek is 6years older than Michelle. In 5 years the sum of their ages will be 58. How old is Derek now
Answer by ikleyn(52807)

.
First equation is


    D - M = 6    (1)


Since in 5 years the sum of their ages will be 58, the sum of their current ages now is  58 - 2*5 = 48.

So, your second equation is


    D + M = 48.


Now add equations (1) and (2). You will get


    2D = 6 + 48 = 54  ====>  D = 54/2 = 27.


Hence,  M = 48 - 27 = 21.


Answer.  D is 27 years old;  M is 21 years old.